For all even integral value of n, lim (x+1)^-n


n to ∞


Exist or not


True or false with full explanation



1
Expert's answer
2022-03-15T12:05:10-0400

"\mathop {\lim }\limits_{n \to \infty } {\left( {x + 1} \right)^{ - n}}"


"= \mathop {\lim }\limits_{n \to \infty } \exp \left( {\ln \left( {{{\left( {x + 1} \right)}^{ - n}}} \right)} \right)"


"= \mathop {\lim }\limits_{n \to \infty } \exp \left( { - n\ln \left( {x + 1} \right)} \right)"


"= \exp \left( { - \ln \left( {x + 1} \right) \cdot \mathop {\lim }\limits_{n \to \infty } n} \right)"


Now for "\ln \left( {x + 1} \right) > 0"


"\mathop {\lim }\limits_{n \to \infty } {\left( {x + 1} \right)^{ - n}}"


"= \exp \left( { - \ln \left( {x + 1} \right) \cdot \infty } \right)"


"= \exp \left( { - \infty } \right)"


"=0"


Hence "\mathop {\lim }\limits_{n \to \infty } {\left( {x + 1} \right)^{ - n}}=0" for "\ln \left( {x + 1} \right) > 0"